A 0.050 M solution of a weak monoprotic acid, HA, has [H3O+] = 3.8 x 10-5. What is the value of Ka for this acid?

2.6 x 10-11

2.9 x 10-8

7.5 x 10-3

7.0 x 10-8

7.0 x 10-7
Answer #1

We know that - ÞH = - dog [H3O+] Given - Ho+1 = 3.8x10-5. .: ÞH = -loq [H30+ ] = - dog (3.8 x 10-5) = 4:42 :ÞH of solution inCriven — сО. ОБО М pH = 4.42 (calculated) Ka=? Accordingly, A pH = /2 bka - Ž soge or, pH + h log c = Ž pka or, 2x (PH + 1/2... The value of Ka for the acid will be 2.9810-8. Am